# FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning

### v4

@@ 1 Introduction
Financial decision–making, such as stochastic modeling, risk assessment, portfolio optimization, and algorithmic trading (Markowitz, 1952; Black and Scholes, 1973a; Rockafellar and Uryasev, 2000; Avellaneda and Stoikov, 2008), demands precise mathematics and domain-specific reasoning Lewkowycz et al. (2022); Wen and Zhang (2025). Recent advances in large foundation models for finance, such as the multimodal FinTral (Bhatia et al., 2024) and language-centric Fin-R1 (Liu et al., 2025), demonstrate progress but still face challenges in interpretability and domain alignment (Nie et al., 2024; Arya.ai, 2024; Lee et al., 2025). Accordingly, these shortcomings motivate stricter control over a model’s intermediate reasoning path, which we explore via prompt design.
@@ 2.2 LLMs in Domain-Specific Financial Reasoning
Methods adapt models via supervised fine-tuning with curated QA/classification datasets (Ma et al., 2023; Harsha et al., 2025) or continued pretraining on domain-specific corpora (Yang et al., 2020; Lee et al., 2024; Bhatia et al., 2024; Ke et al., 2025; Liu et al., 2025). While effective, these approaches rely on labeled data and general language modeling objectives, lacking structuring of financial reasoning, thus limiting interpretability and alignment with expert logic.
@@ 5.1 Financial Reasoning Performance
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Qwen3-8B
Gemma-3-12B
IT
Fin-R1
7B
DianJin-R1
7B
Fin-o1
8B
@@
FinCoT also helps specialized models like Fin-R1, confirming that blueprint prompting provides complementary gains beyond fine-tuning. However, for models with strong built-in reasoning, such as DianJin-R1-7B and Fin-o1-8B, FinCoT offers limited improvement or slight degradation—likely due to conflicts between external scaffolds and internal reasoning routines. These outcomes suggest diminishing returns for CoT prompting when domain alignment and reasoning are already deeply encoded.
@@ Appendix A Expert Reasoning Blueprints
Step-2: Identify Topic (A1) – Microeconomics (A2): Focus on supply & demand mechanisms and market structures such as perfect competition, monopoly, oligopoly, and monopolistic competition (Investopedia, 2025a; 300Hours, 2025a). – Macroeconomics (A3): Consider aggregate demand–aggregate supply analysis, phases of the business cycle (expansion, peak, contraction, trough), and growth models (Solow, endogenous growth) (Investopedia, 2025b; CFA Institute, 2024). – Currency & Regulation (A4): Examine exchange-rate regimes (floating vs. pegged), trade balances, capital-flow impacts, and relevant government policies (Investopedia, 2025, 2025b).
Step-3: Model Selection or Strategy Mapping (B1–B3) – Micro Models (B1): Choose elasticity calculations and cost-curve analysis (marginal/average cost, shutdown point) for supply–demand or firm-behavior questions (Investopedia, 2025e). – Macro Frameworks (B2): Apply AD–AS curves, Phillips-curve trade-offs, or business-cycle indicators to frame policy or growth analysis (Investopedia, 2025, 2025c). – FX & Regulation (B3): Use exchange-rate determination models, balance-of-payments analysis, or regulatory impact frameworks for currency/trade questions (Investopedia, 2025a).
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $\text{GDP}=C+I+G+(X-M)$ ), and verify assumptions (Investopedia, 2025c). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary-policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025b).
Step-5: Solution Development (E) – Structured Response (E1): State the core economic insight first, then provide the step-by-step rationale linking theory to the question context (Kaplan Schweser, 2025). – Alternative Scenarios (E2): Where relevant, outline best-case, base-case, and worst-case scenarios or show how supply–demand curves shift under different assumptions (iPassFinanceExams, 2025).
@@
300Hours CFA Level 1 Economics Cheat Sheet (300Hours, 2025a), UWorld Finance’s CFA® Economics: Syllabus & Sample Questions (UWorld Finance, 2025a), Kaplan Schweser’s Level I Economics tips (Kaplan Schweser, 2025), Efficient Learning’s CFA Economics overview (Efficient Learning, 2025), iPass Finance Exams’ study guide (iPassFinanceExams, 2025), AnalystPrep’s essential review summary (AnalystPrep, 2025), and key Investopedia articles on supply and demand (Investopedia, 2025a), GDP (Investopedia, 2025c), and business cycles (Investopedia, 2025b) as well as Prof. Brian Gordon’s CFA Exam Level 1 Economics video (Gordon, 2020a).
@@
Step-1: Purpose and Scope – Define the investment objective–income generation, capital preservation, hedging, or total return and establish portfolio constraints and benchmarks such as target yield, duration limits, credit quality floors, or sector allocation guidelines (Investopedia, 2025b).
Step-2: Analyze Macro Conditions – Examine current and forecast interest rate paths, since rising rates erode bond prices and falling rates support them (Investopedia, 2025); monitor inflation indicators (CPI, PPI) to gauge real yield trends (Investopedia, 2025a); and assess yield-curve shapes (normal, inverted, flat) for economic turning points and yield-curve trade opportunities (Investopedia, 2025b).
Step-3: Assess Bond Features – Identify bond type government, corporate, municipal, structured products (ABS/MBS) and note any embedded options (callable, putable, convertible) (Investopedia, 2025); review coupon structure (fixed vs. floating), payment frequency, and maturity to understand cash-flow timing and reinvestment risk (Investopedia, 2025c).
Step-4. Risk and Yield Analysis – Calculate duration to estimate price sensitivity to yield changes (Investopedia, 2025e) and convexity for non-linear price effects (Investopedia, 2025b); analyze credit spreads over benchmarks to gauge default and liquidity risk (Investopedia, 2025d); and stress-test the portfolio under parallel shifts, steepeners, and flatteners to assess P&L impacts.
@@
Step-6: Review Performance – Track total returns (price changes plus coupon income) against benchmarks like the Bloomberg US Aggregate Bond Index (Bloomberg, 2025); perform attribution analysis to decompose yield carry, curve roll-down, and spread effects; and revisit assumptions and rebalance when market conditions or issuer fundamentals change (Investopedia, 2025a).
@@
CFA Program Curriculum for Fixed Income (CFA Institute, 2025e), Fabozzi’s Bond Markets, Analysis, and Strategies (Fabozzi, 2012), Tuckman & Serrat’s Fixed Income Securities (Tuckman and Serrat, 2011), Jarrow & Turnbull’s credit-risk derivatives pricing (Jarrow and Turnbull, 1995), Basel Committee papers on credit risk, Investopedia articles on fixed-income concepts (Investopedia, 2025b, b, e), Reuters coverage of convexity risk, and Dichev’s balance sheet model for distress prediction (Dichev, 2017).
@@
Step-1: Articulating Purpose and Context (A) Define the research question time value of money calculations, probability distributions, hypothesis testing, regression analysis, or portfolio statistics–and establish the CFA application context (market efficiency, risk estimation, cash-flow forecasting) (CFA Institute, 2025h).
Step-2: Collecting Input Data (B) Gather historical returns, economic indicators, financial statements, and market data from reputable sources; ensure relevance by matching data to the chosen objective (e.g., interest rates for TVM, volatility for risk models) (Investopedia, 2025f).
@@
Step-4: Selecting Quantitative Models and Tools (D) Choose appropriate models–ARIMA for time series, linear/multivariate regression, probability distributions, or Monte Carlo simulation–and leverage CFA-recommended software or spreadsheet tools (Damodaran, 2012; Investopedia, 2025g).
@@
Step-1: Objective Setting (A) Define your investment objectives–capital appreciation, dividend income, or total return in line with your risk tolerance and investment horizon; consider constraints such as liquidity needs, tax implications, regulatory requirements, and any specific mandates (CFA Institute, 2025b).
Step-2: Market and Sector Insights (B) Assess macro indicators (GDP growth, interest rates, inflation) to gauge the overall market environment and identify sectors poised for growth or decline based on economic trends, technological shifts, and consumer behavior (Investopedia, 2025a).
Step-3: Industry Competitive Analysis (C) Apply Porter’s Five Forces to evaluate industry attractiveness–competitive rivalry, threat of new entrants, bargaining power of suppliers and buyers, and substitute threats–and assess each firm’s market share and competitive moat (Investopedia, 2025d).
Step-4: Company Review (D) Examine financial statements (income statement, balance sheet, cash flows) to measure profitability, liquidity, and stability; evaluate management’s track record and strategic vision; and review corporate governance structures to ensure alignment with shareholder interests (Investopedia, 2025d; Bodie et al., 2017).
@@
CFA Program Curriculum’s Equity Investments module (CFA Institute, 2025b), Investopedia’s guides on fundamental analysis (Investopedia, 2025a), Porter’s Five Forces stock analysis (Investopedia, 2025d), and reading financial reports (Investopedia, 2025d), 300Hours’ CFA Level 1 Equity Cheat Sheet (300Hours, 2025b), Bodie, Kane & Marcus’s Investments (Bodie et al., 2017), Pinto et al.’s Equity Asset Valuation (Pinto et al., 2015), and CFA Level I Equity video lectures by Prof. Brian Gordon (Gordon, 2020b, c).
@@
  • Black-Scholes model for European options (Black and Scholes, 1973b).
@@
Based on Hull’s comprehensive treatment of markets and pricing models (Hull, 2017), the CFA Institute Level II Derivatives readings (CFA Institute, 2025f), Black & Scholes’s seminal option pricing model (Black and Scholes, 1973b), Jarrow & Turnbull’s practical engineering perspective (Jarrow and Turnbull, 1996), Tavella & Randall’s numerical finite-difference techniques (Tavella and Randall, 2000), and the Basel Committee’s OTC derivatives reforms for regulatory context (Board, 2025).
@@
Gather primary financial statements: income statement, balance sheet, and cash flow statement. Supplement this with industry benchmarks and macroeconomic data. Ensure the quality, accuracy, and completeness of all collected data (Investopedia, 2025a).
@@
CFA Program Curriculum’s Financial Reporting and Analysis readings covering ratio analysis, cash flow analysis, and IFRS/GAAP standards (CFA Institute, 2025d) alongside Investopedia’s overview of financial statement components (Investopedia, 2025a), Paul R. Brown’s strategic perspective on statement analysis and valuation (Stickney et al., 2007), and Palepu & Healy’s MBA-level treatment of business analysis and valuation using financial statements (Palepu et al., 2013).
@@
Step-1: Corporate Issuer Overview – Begin with a high-level understanding of the firm’s business model, market positioning, and strategic objectives. This foundational context is essential for both equity and fixed income analysis (CFA Institute, 2025c).
@@
Step-3: Financial Statement Analysis and Key Metrics – Analyze income statement, balance sheet, and cash flow data. Focus on metrics like revenue growth, operating margin, return on equity, and leverage. This step reveals the firm’s financial health and operational efficiency (Penman, 2012; CFA Institute, 2025c).
@@
Step-6: Corporate Governance and Leadership Quality – Assess governance practices such as board independence, shareholder rights, and disclosure quality. Evaluate the management team’s execution track record and alignment with shareholder interests (CFA Institute, 2025c).
@@
CFA Program Curriculum’s Equity Investments and Fixed Income readings–which cover firm analysis, industry evaluation, and credit assessment frameworks (CFA Institute, 2025c)–along with Penman’s Financial Statement Analysis and Security Valuation for accounting-to-valuation linkages (Penman, 2012), and Fabozzi’s Bond Markets, Analysis, and Strategies for credit risk and corporate debt issuance insights (Fabozzi, 2012).
@@ D.1 Average Input Tokens
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Gemma-3-12B
IT
Qwen3-8B
Fin-R1
DianJin-R1
7B
Fin-o1-8B

### v5

@@ 1 Introduction
Financial decision–making, such as stochastic modeling, risk assessment, portfolio optimization, and algorithmic trading (Markowitz, 1952; Black and Scholes, 1973; Rockafellar and Uryasev, 2000; Avellaneda and Stoikov, 2008), demands precise mathematics and domain-specific reasoning Lewkowycz et al. (2022); Wen and Zhang (2025). Recent advances in large foundation models for finance, such as the multimodal FinTral (Bhatia et al., 2024) and language-centric Fin-R1 (Liu et al., 2025), demonstrate progress but still face challenges in interpretability and domain alignment (Nie et al., 2024; Arya.ai, 2024; Lee et al., 2025). Accordingly, these shortcomings motivate stricter control over a model’s intermediate reasoning path, which we explore via prompt design.
@@ 2.2 LLMs in Domain-Specific Financial Reasoning
Methods adapt models via supervised fine-tuning with curated QA/classification datasets (Chen et al., 2022; Harsha et al., 2025) or continued pretraining on domain-specific corpora (Yang et al., 2020; Lee et al., 2024; Bhatia et al., 2024; Ke et al., 2025; Liu et al., 2025). While effective, these approaches rely on labeled data and general language modeling objectives, lacking structuring of financial reasoning, thus limiting interpretability and alignment with expert logic.
@@ 5.1 Financial Reasoning Performance
Qwen2.5-7B Qwen2.5-7B Instruct Qwen3-8B Base Qwen3-8B Qwen3-8B (Thinker) Gemma-3-12B IT Fin-R1 7B DianJin-R1 7B Fin-o1 8B
@@
FinCoT also helps specialized models like Fin-R1, confirming that blueprint prompting provides complementary gains beyond fine-tuning. However, for models with strong built-in reasoning, such as DianJin‑R1‑7B and Fin‑o1‑8B, FinCoT offers limited improvement or slight degradation—likely due to conflicts between external scaffolds and internal reasoning routines. These outcomes suggest diminishing returns for CoT prompting when domain alignment and reasoning are already deeply encoded.
@@ Appendix A Expert Reasoning Blueprints
Step-2: Identify Topic (A1) – Microeconomics (A2): Focus on supply & demand mechanisms and market structures such as perfect competition, monopoly, oligopoly, and monopolistic competition (Investopedia, 2025r; 300Hours, 2025a). – Macroeconomics (A3): Consider aggregate demand–aggregate supply analysis, phases of the business cycle (expansion, peak, contraction, trough), and growth models (Solow, endogenous growth) (Investopedia, 2025d; CFA Institute, 2024). – Currency & Regulation (A4): Examine exchange-rate regimes (floating vs. pegged), trade balances, capital-flow impacts, and relevant government policies (Investopedia, 2025j, n).
Step-3: Model Selection or Strategy Mapping (B1–B3) – Micro Models (B1): Choose elasticity calculations and cost‐curve analysis (marginal/average cost, shutdown point) for supply–demand or firm‐behavior questions (Investopedia, 2025v). – Macro Frameworks (B2): Apply AD–AS curves, Phillips‐curve trade‐offs, or business‐cycle indicators to frame policy or growth analysis (Investopedia, 2025a, t). – FX & Regulation (B3): Use exchange‐rate determination models, balance‐of‐payments analysis, or regulatory impact frameworks for currency/trade questions (Investopedia, 2025c).
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $\text{GDP}=C+I+G+(X-M)$ ), and verify assumptions (Investopedia, 2025o). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary‐policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025s).
Step-5: Solution Development (E) – Structured Response (E1): State the core economic insight first, then provide the step‐by‐step rationale linking theory to the question context (Kaplan Schweser, 2025). – Alternative Scenarios (E2): Where relevant, outline best‐case, base‐case, and worst‐case scenarios or show how supply–demand curves shift under different assumptions (iPassFinanceExams, 2025).
@@
300Hours CFA Level 1 Economics Cheat Sheet (300Hours, 2025a), UWorld Finance’s CFA® Economics: Syllabus & Sample Questions (UWorld Finance, 2025a), Kaplan Schweser’s Level I Economics tips (Kaplan Schweser, 2025), Efficient Learning’s CFA Economics overview (Efficient Learning, 2025), iPass Finance Exams’ study guide (iPassFinanceExams, 2025), AnalystPrep’s essential review summary (AnalystPrep, 2025), and key Investopedia articles on supply and demand (Investopedia, 2025r), GDP (Investopedia, 2025o), and business cycles (Investopedia, 2025d) as well as Prof. Brian Gordon’s CFA Exam Level 1 Economics video (Gordon, 2020).
@@
Step-1: Purpose and Scope – Define the investment objective–income generation, capital preservation, hedging, or total return and establish portfolio constraints and benchmarks such as target yield, duration limits, credit quality floors, or sector allocation guidelines (Investopedia, 2025l).
Step-2: Analyze Macro Conditions – Examine current and forecast interest rate paths, since rising rates erode bond prices and falling rates support them (Investopedia, 2025q); monitor inflation indicators (CPI, PPI) to gauge real yield trends (Investopedia, 2025e); and assess yield-curve shapes (normal, inverted, flat) for economic turning points and yield-curve trade opportunities (Investopedia, 2025z).
Step-3: Assess Bond Features – Identify bond type government, corporate, municipal, structured products (ABS/MBS) and note any embedded options (callable, putable, convertible) (Investopedia, 2025b); review coupon structure (fixed vs. floating), payment frequency, and maturity to understand cash-flow timing and reinvestment risk (Investopedia, 2025g).
Step-4. Risk and Yield Analysis – Calculate duration to estimate price sensitivity to yield changes (Investopedia, 2025i) and convexity for non-linear price effects (Investopedia, 2025f); analyze credit spreads over benchmarks to gauge default and liquidity risk (Investopedia, 2025h); and stress-test the portfolio under parallel shifts, steepeners, and flatteners to assess P&L impacts.
@@
Step-6: Review Performance – Track total returns (price changes plus coupon income) against benchmarks like the Bloomberg US Aggregate Bond Index (Bloomberg, 2025); perform attribution analysis to decompose yield carry, curve roll-down, and spread effects; and revisit assumptions and rebalance when market conditions or issuer fundamentals change (Investopedia, 2025y).
@@
CFA Program Curriculum for Fixed Income (CFA Institute, 2025e), Fabozzi’s Bond Markets, Analysis, and Strategies (Fabozzi, 2012), Tuckman & Serrat’s Fixed Income Securities (Tuckman and Serrat, 2011), Jarrow & Turnbull’s credit‐risk derivatives pricing (Jarrow and Turnbull, 1995), Basel Committee papers on credit risk, Investopedia articles on fixed‐income concepts (Investopedia, 2025l, f, i), Reuters coverage of convexity risk, and Dichev’s balance sheet model for distress prediction (Dichev, 2008).
@@
Step-1: Articulating Purpose and Context (A) Define the research question time value of money calculations, probability distributions, hypothesis testing, regression analysis, or portfolio statistics–and establish the CFA application context (market efficiency, risk estimation, cash‐flow forecasting) (CFA Institute, 2025h).
Step-2: Collecting Input Data (B) Gather historical returns, economic indicators, financial statements, and market data from reputable sources; ensure relevance by matching data to the chosen objective (e.g., interest rates for TVM, volatility for risk models) (Investopedia, 2025w).
@@
Step-4: Selecting Quantitative Models and Tools (D) Choose appropriate models–ARIMA for time series, linear/multivariate regression, probability distributions, or Monte Carlo simulation–and leverage CFA-recommended software or spreadsheet tools (Damodaran, 2012; Investopedia, 2025x).
@@
Step-1: Objective Setting (A) Define your investment objectives–capital appreciation, dividend income, or total return in line with your risk tolerance and investment horizon; consider constraints such as liquidity needs, tax implications, regulatory requirements, and any specific mandates (CFA Institute, 2025c).
Step-2: Market and Sector Insights (B) Assess macro indicators (GDP growth, interest rates, inflation) to gauge the overall market environment and identify sectors poised for growth or decline based on economic trends, technological shifts, and consumer behavior (Investopedia, 2025m).
Step-3: Industry Competitive Analysis (C) Apply Porter’s Five Forces to evaluate industry attractiveness–competitive rivalry, threat of new entrants, bargaining power of suppliers and buyers, and substitute threats–and assess each firm’s market share and competitive moat (Investopedia, 2025u).
Step-4: Company Review (D) Examine financial statements (income statement, balance sheet, cash flows) to measure profitability, liquidity, and stability; evaluate management’s track record and strategic vision; and review corporate governance structures to ensure alignment with shareholder interests (Investopedia, 2025p; Bodie et al., 2017).
@@
CFA Program Curriculum’s Equity Investments module (CFA Institute, 2025c), Investopedia’s guides on fundamental analysis (Investopedia, 2025m), Porter’s Five Forces stock analysis (Investopedia, 2025u), and reading financial reports (Investopedia, 2025p), 300Hours’ CFA Level 1 Equity Cheat Sheet (300Hours, 2025b), Bodie, Kane & Marcus’s Investments (Bodie et al., 2017), Pinto et al.’s Equity Asset Valuation (Pinto et al., 2015), and CFA Level I Equity video lectures by Prof. Brian Gordon (PrepNuggets, 2020; CFA Society Nigeria, 2024).
@@
  • Black-Scholes model for European options (Black and Scholes, 1973).
@@
Based on Hull’s comprehensive treatment of markets and pricing models (Hull, 2017), the CFA Institute Level II Derivatives readings (CFA Institute, 2025f), Black & Scholes’s seminal option pricing model (Black and Scholes, 1973), Jarrow & Turnbull’s practical engineering perspective (Jarrow and Turnbull, 1996), Tavella & Randall’s numerical finite‐difference techniques (Tavella and Randall, 2000), and the Basel Committee’s OTC derivatives reforms for regulatory context (Board, 2025).
@@
Gather primary financial statements: income statement, balance sheet, and cash flow statement. Supplement this with industry benchmarks and macroeconomic data. Ensure the quality, accuracy, and completeness of all collected data (Investopedia, 2025k).
@@
CFA Program Curriculum’s Financial Reporting and Analysis readings covering ratio analysis, cash flow analysis, and IFRS/GAAP standards (CFA Institute, 2025d) alongside Investopedia’s overview of financial statement components (Investopedia, 2025k), Paul R. Brown’s strategic perspective on statement analysis and valuation (Stickney et al., 2007), and Palepu & Healy’s MBA‐level treatment of business analysis and valuation using financial statements (Palepu et al., 2013).
@@
Step-1: Corporate Issuer Overview – Begin with a high-level understanding of the firm’s business model, market positioning, and strategic objectives. This foundational context is essential for both equity and fixed income analysis (CFA Institute, 2025b).
@@
Step-3: Financial Statement Analysis and Key Metrics – Analyze income statement, balance sheet, and cash flow data. Focus on metrics like revenue growth, operating margin, return on equity, and leverage. This step reveals the firm’s financial health and operational efficiency (Penman, 2012; CFA Institute, 2025b).
@@
Step-6: Corporate Governance and Leadership Quality – Assess governance practices such as board independence, shareholder rights, and disclosure quality. Evaluate the management team’s execution track record and alignment with shareholder interests (CFA Institute, 2025b).
@@
CFA Program Curriculum’s Equity Investments and Fixed Income readings–which cover firm analysis, industry evaluation, and credit assessment frameworks (CFA Institute, 2025b)–along with Penman’s Financial Statement Analysis and Security Valuation for accounting-to-valuation linkages (Penman, 2012), and Fabozzi’s Bond Markets, Analysis, and Strategies for credit risk and corporate debt issuance insights (Fabozzi, 2012).
@@ D.1 Average Input Tokens
Qwen2.5-7B Qwen2.5-7B Instruct Qwen3-8B Base Qwen3-8B Gemma-3-12B IT Qwen3-8B (Thinker) Fin-R1 DianJin-R1 7B Fin-o1-8B